### v4

# Ferromagnetic CrBr3-Induced Graphene Spintronics

Abstract
[…]
Introduction–Spin valves are crucial in spintronics, composed of layered ferromagnetic materials as conducting electrodes RevModPhys.81.109; PhysRevB.43.1297. They are integral to modern electronics, leveraging giant magnetoresistance (MR) to switch between high and low resistance states monsma1995perpendicular; park2011aspinvalve; grunberg2008from; acs.nanolett.6b03052. Understanding two-dimensional (2D) van der Waals (vdWs) systems has laid a robust foundation, offering finely tunable electronic properties in heterostructures adma.202209113; acs.jpclett.3c03000; sciadv.abn1401; Flokstra2023; PhysRevLett.126.056803; PhysRevLett.128.167701 with materials like h-BN, TMDCs (e.g., MoS2), graphene, and phosphorene Geim2013; Novoselovaac9439; PhysRevLett.114.066803; nat2019years; han2014graphene. In vdW spin valves, the tunable spin proximity effect (SPE) can induce a metal-insulator transition in graphene under strong magnetic fields, potentially leading to Quantum Hall ferromagnetism PhysRevLett.95.146802; PhysRevB.82.161414; PhysRevLett.108.056802; PhysRevB.83.155447; PhysRevB.85.115439; acs.nanolett.3c02489; acs.nanolett.2c03113. Advances with monolayers of ferromagnetic insulator CrBr3 in vdWs devices have opened avenues for exploring spin-dependent transport phenomena Tombros2007; Song1214.
Recently, a van der Waals (vdW) spin valve was demonstrated using bilayer graphene (BLG) between CrI3 ferromagnetic insulators PhysRevLett.121.067701. In our study, we aim to explore a similar configuration but replacing CrI3 with CrBr3, which has a slightly lower Curie temperature (TC) of 33K PhysRevB.56.719 compared to CrI3 (TC of 45K) Huang2017; PhysRevLett.121.067701. CrBr3 is notable for its heavy Cr atoms, suggesting potential for strong spin-orbit coupling (SOC) effects in graphene PhysRevB.84.195444; PhysRevLett.109.055502; acsami.4c01034. Its hexagonal crystal structure aligns well with graphene’s lattice, with a modest lattice mismatch of approximately 4.5 $\%$. This structural compatibility makes CrBr3 promising for vdW heterostructures. The layered architecture of CrBr3, bound by vdW forces, allows easy exfoliation into 2D monolayers while maintaining ferromagnetic insulating properties PhysRevLett.110.046603; Hallal_2017; acs.nanolett.9b00553, enhancing the versatility of vdW spin valves. In CrBr3 bilayers, weak interlayer coupling is conducive to controlling layer magnetizations independently, crucial for spintronic applications PhysRevB.84.195444; PhysRevLett.109.055502. This exploration opens new avenues for studying spin-dependent phenomena in 2D vdW heterostructures.
Our study focuses on bilayer graphene, diverging from previous lateral spin valves that utilized monolayer graphene Tombros2007; han2014graphene. Monolayer graphene has limited contact area with ferromagnetic electrodes, unlike spin filter tunnel junctions Klein1218; Song1214; Wang2018. In contrast, our proposed van der Waals (vdW) spin valve integrates bilayer graphene between insulating magnetic layers PhysRevLett.121.067701; PhysRevLett.61.2472; PhysRevB.39.4828. This design offers advantages such as increased contact area between the graphene and magnetic materials, potentially enhancing spin transport properties. Moreover, bilayer graphene’s unique electronic structure and tunable properties enable precise control over spin-related phenomena. These attributes highlight the novelty and potential effectiveness of our vdW spin valve architecture.
Figure 1: (a) Atomic structure of a CrBr3-BLG-CrBr3 heterostructure unit cell. Inter-planner spacing notation of the structure in (b) and molecular surface formation of the heterostructure showing distribution (c). Atomic structure of a CrBr3-graphene bilayer-CrBr3 heterostructure supercell of 5*5*1 (d). Polyhedra presentation of the same heterostructure showing the electron bond and distribution (e).
Computational Details–The geometry optimization and electronic structure calculations of the vdW heterostructure were conducted using the projector augmented wave (PAW) formalism PhysRevB.50.17953 within density functional theory (DFT) PhysRevB.54.11169. The Perdew-Burke-Ernzerhof (PBE) generalized-gradient approximation (GGA) was employed to describe the exchange and correlation functional PhysRevLett.77.3865. Specifically for the Cr 3d electrons, the GGA+U method was applied, with the effective on-site Coulomb interaction U set to 3.0 eV and exchange interaction J to 0.9 eV PhysRevB.52.R5467. For the monolayer CrBr3, known for its accuracy but higher computational expense, the HSE06 hybrid functional 10.1063/1.1564060 was adopted to achieve optimized geometry structures and electronic states. We design this heterostructure with a bilayer graphene (BLG) sheet surrounded by two monolayers of CrBr3. For the bilayer structure, we used a unit cell composed of 4 layers: CrBr3/BLG/CrBr3. In the superlattice arrangement, we adjusted the in-plane lattice parameter of graphene while varying the $c$ lattice parameter to optimize the interlayer distance (2.34 Å between CrBr3/BLG and 1.8 Å between BLG layers) and minimize the total energy. Each unit cell in the superlattice setup consisted of a $\times3$$\times$1 supercell, containing 14 carbon atoms per graphene layer, along with 2 chromium and 6 bromine atoms per CrBr3 layer. Consequently, the sandwiched graphene bilayer structure encompassed a total of 44 atoms. Our calculations revealed that the spin of Cr atoms is $S=3/2$, residing within the t2gt{2g} bands. A plane-wave cutoff energy of 550 eV and a vacuum space larger than 22 Å were implemented to prevent interactions between adjacent slabs. Grids of $\times9$$\times$1 and $\times27$$\times$1 Monkhorst-Pack k-point meshes were used for the vdW heterostructures and pure monolayer CrBr3, respectively, to integrate over the first Brillouin zone. Here, we consider both the parallel configurations showing ferromagnetic (FM) and antiferromagentic (AFM) orientations. The vdW correction was incorporated using the Grimme (DFT-D2) method 10.1002.20495 in the heterostructure calculations. Electronic band structures were also computed using the Quantum ESPRESSO suite Giannozzi_2009 to support and validate the computational findings.
Figure 2: DFT band structure of the heterostructure with inset showing finite gap opening of 56.87 meV. Blue colour and red colour dotted lines present both majority and minority spins of both systems, respectively along highly symmetric k-points.
Figure 3: DFT calculation results in bilayer graphene sandwiched by CrBr3. (a) The energy difference between ferromagnetic and antiferromagnetic states. (b) Spin-dependent gap as a function of U in presence of field.
Results and Discussions–The atomic configuration is depicted in Fig.1(a-c), and the supercell architecture is illustrated in Fig.1(d-e). The interlayer distance between the graphene sheet and CrBr3 is 2.34 Å, while the interplanar distance between graphene sheets is fixed at 1.80 Å, characterized by van der Waals-type interlayer interactions. The total optimized energy is simulated as a function of interplanar distance to determine the optimal spacing for effective proximity coupling between the graphene sheet and CrBr3, incorporating dispersion corrections monsma1995perpendicular.
Dispersion correction has been integrated to accurately assess the bonding compatibility between the CrBr3 layer and the graphene sheet, ensuring precise determination of the total energy relative to the interplanar spacing. In constructing graphene-on-metal systems as heterostructures, it is crucial to comprehend and minimize junction resistance originating from the metal-graphene interface during geometry optimization, alongside implementing dispersion corrections for van der Waals interacting systems park2011aspinvalve; grunberg2008from; han2014graphene.
Figure 4: DFT local potential distribution of the CrBr3-BLG-CrBr3 heterostructure with contribution from both CrBr3 monolayer and graphene bilayer along z-direction.
The electronic band structure of the CrBr3/BLG/CrBr3 heterostructure at zero biasing is depicted in Fig. 2 under a ferromagnetic state. Analysis of the band structure reveals prominent graphene energy bands near the Fermi level, with an observed gap of 56.87 meV (inset of Fig. 2. This characteristic illustrates well-organized energy levels at the K point, accounting for spin degeneracy. Consequently, the proximity magnetism exerted by the CrBr3 layers on the BLG system manifests a discernible AFM ground state (depicted in Fig. 2, unlike FM state. To support the electronic band structure analysis, the band structure of our model vdW spin valve is depicted in Figure 2 alongside the $e_{g}$ bands. In this scenario, the shape of the graphene bilayer bands remains unchanged in the minority spin channel, while significant hybridization introduces a gap in the majority spin channel just above the Fermi energy, positioned below the k point. In contrast, both spin channels of the graphene bilayer undergo hybridization with the narrow $e_{g}$ bands of CrBr3. This analysis suggests a notably higher in-plane conductance in the AFM alignment based solely on this spintronics perspective.
We assume that the magnetizations of both the top and bottom layers are aligned along the same axis during the calculations. The magnetization of the bottom layer is assumed to be fixed, leading to a constant spin-dependent potential of $\Sigma\delta$, where $\Sigma$ represents the spin projection along this axis. For the top layer, the spin-dependent potential is given by $\epsilon\Sigma\delta$, where $\epsilon=\pm 1$ indicates the relative orientation of the top layer’s magnetization with respect to the bottom layer. Specifically, $\epsilon=+1$ corresponds to a parallel FM alignment, while $\epsilon=-1$ corresponds to an antiparallel AFM alignment. In this case, the Bloch Hamiltonian for spin $\Sigma$ states is expressed as follows,:
${\cal H}_{\Sigma}(\vec{k})=\left(\begin{array}[]{cccc}\epsilon\Sigma\frac{\delta}{2}&f(\vec{k})&0&0\\ f^{*}(\vec{k})&\epsilon\Sigma\frac{\delta}{2}&\Gamma&0\\ 0&\Gamma&\Sigma\frac{\delta}{2}&f(\vec{k})\\ 0&0&f(\vec{k})^{*}&\Sigma\frac{\delta}{2}\end{array}\right)$ (5)
Here, $f(\vec{k})=t\left(1+e^{i\vec{k}\cdot\vec{a}_{1}}+e^{i\vec{k}\cdot\vec{a}_{2}}\right)$ and $\Gamma$ denote the intralayer and interlayer hopping matrix elements, respectively. The spin-resolved energy bands near the Dirac K point are illustrated in Fig. 2 for the spin valve architecture. In the FM alignment ($\epsilon=+1$), the BLG system exhibits spin-split bands and remains in a metallic state, meaning it has a finite density of states at the Fermi energy. Conversely, in the AFM alignment ($\epsilon=-1$), a band gap (56.87 meV) opens at the Dirac K point. Thus, the BLG spin valve can act as a conductor in the FM state or as a gapped insulator at the Dirac K point in the AFM state.
These results highlight the potential of using a vdW magnetic system to explore quantum Hall effects, diverging from conventional ferromagnetic layer systems. Examining the band structure reveals fluctuations in the split-off energy gap between the ferromagnetic and antiferromagnetic states (illustrated in Fig.3(a)), along with the spin-dependent gap as a function of U across bias voltages from 0 to ±5 V/nm (shown in Fig.3(b)). This analysis demonstrates pronounced non-linearity under biasing conditions, indicative of interlayer charge polarization. Moreover, the mixed (both linear and nonlinear) behavior prompts further investigation into charge transmission spectra to understand the proximity effect in heterostructures and the role of interlayer polarization for potential device applications (as depicted in the inset heterostructure in Fig. 2). We now explore the origin of gap formation in the graphene bilayer within an antiferromagnetic aligned vdW spin valve. In this setup, the Hamiltonian for each spin channel mirrors the model of a graphene bilayer under an off-plane electric field, which is known to induce a band gap in the structure Ohta951; PhysRevB.74.161403. Notably, within the spin valve configuration, the effective electric field exhibits opposite signs for different spin orientations, $E_{\rm eff}\propto\sigma\Delta$. The spin projection of the antiferromagnetic bands in the top and bottom layers, as depicted in Fig. 3(b), distinctly manifests a spin dipole effect: for a given spin direction, there arises a charge imbalance induced by the exchange interaction with the magnetic layers, precisely offset by the opposite spin direction.
We explore the energy states around the Fermi energy ($E_{F}$) of a graphene bilayer, where the valence and conduction parabolic bands converge. This region shows a slight band splitting, indicative of parabolic electrons and holes in the BLG structure, with a split of 56.87 meV observed for the AFM case (as depicted in the inset of Fig. 2). Similarly, there is a distinct spin splitting of the bands near the k point in this configuration, with the same magnitude as the AFM gap. This consistency supports the DFT findings for bilayer graphene sandwiched between two CrBr3 layers, consistent with theoretical predictions. However, due to electron transfer from graphene to CrBr3, the Fermi energy does not precisely align at the k point. It’s noteworthy that despite this charge transfer, CrBr3 exhibits very limited in-plane conductance due to the minimal dispersion of occupied states, as evidenced by the local potential distribution plot (see Fig. 4), showing the contributions of BLG and CrBr3 layers.
Summary–We have proposed a novel spin valve design where a non-magnetic 2D crystal is sandwiched between two ferromagnetic insulators. Its in-plane conductance, mediated through nanoscale transport, is controlled by the spin proximity effect. We anticipate that our work will inspire experimentalists to explore van der Waals (vdW) in-plane spin valves further. This includes investigating other materials for the non-magnetic layers, such as superconductors, as well as alternative magnetic layers, like bulk ferromagnets, for which the proposed spin valve effect should also be applicable.
keywords– vdW heterostructure, ${CrBr}_{3}$, Spin proximity effect, Spintronics, Density functional theory